OpenStudy (anonymous):

Evaluate the following expression 5^-3 3^-2

OpenStudy (anonymous):

1/25 and 1/9 I think

OpenStudy (johnweldon1993):

\[\large 5^{-3} = \frac{1}{5^3} = \frac{1}{125}\]

OpenStudy (johnweldon1993):

same with 3^-2 \[\large 3^{-2} = \frac{1}{3^2} = \frac{1}{9}\]

OpenStudy (anonymous):

Yea sorry I meant 1/125 not 1/25